# Improved Contact Graph Routing in Delay Tolerant Networks with Capacity and Buffer Constraints

### v1

@@ IV-C Booking resources for forwarding safety margin
With the proposed solution only source nodes keep track of the buffer and contact capacity occupancy or status. Due to this feature, they can avoid overflows and overbooking. Nevertheless, due to some unpredictable disruptions, like antenna pointing errors or node power outages [19], intermediate nodes may have to perform routing during forwarding time. Since intermediate nodes may not have access to buffer and capacity status, their routing decision may cause overflows or overbooking. For this reason, we propose to let parts of the original contact and buffer resources unused by source nodes. These unused portions are referred to as safety margins and are utilized by intermediate nodes to reroute bundles without resulting in overflows or overbooking. The quantity of resources to be allocated through this method can be determined as a function of the network’s disruption level.
These safety margins can be implemented as follows. First, regarding the contact-capacity management, two CPs are created. Contact splitting, for the slots that should be kept as margin, is performed. The safety margin is therefore removed from the original CP. The obtained CP is provided to and used by the source nodes. The safety margin part (contacts removed from the original CP) is forwarded to intermediate nodes that may perform routing in unpredictable situations.
In terms of buffer management, source nodes utilize a fraction of the buffer capacity, leaving the rest available for intermediate nodes. This solution therefore enhances robustness to potential unexpected events, despite the poor knowledge of the current network status by intermediate nodes.
@@ IV-D Information sharing in the network
As mentioned above, the modified network information after contact splitting (except the bundle-specific CP modifications) and the modified forecast buffer tables (after allocating a route for a bundle) are to be shared to all source nodes performing a route search. This is similar to the sharing of the up-to-date CP to the source nodes in standard SR-CGR.
The safety margin resources are shared to all nodes in the network at a low frequency.
@@ V Benchmark: source-routing CGR
As mentioned in the first part of the manuscript, network issues are handled after the route search in standard SR-CGR. While retransmission mechanism do exist for long round-trip times communication, such as Licklider Transmission Protocol (LTP) [20], retransmission mechanism in case of failure in standard SR-CGR is not well described in the literature. As there are many manners to address these issues, we describe in details the assumptions made for the considered SR-CGR benchmark. The proposed algorithm is compared to this benchmark in Section VII, where the simulation results are presented.
To begin with, both in the benchmark and in our algorithm, when a bundle is being transmitted over a contact, it is assumed that all the link is occupied for the time required to transmit the bundle. In other words, the whole link capacity is used (i.e., Time Division Multiplexing (TDM) approach). The initial CP is constructed at the source and is broadcasted to all nodes in the network.
Then, the source node performs route search using Yen’s algorithm with Dijkstra searches. A list of the $k$ (Yen’s constant) shortest routes is generated. This list is kept until a change occurs in the CP, such as the end of a contact. In this case, the route list is discarded and a new list is generated, as in route list pruning [12]. A list of candidate routes is constructed as explained in Section II-B3. When no candidate route is found in this list, the bundle is delayed and the route search is performed again after one time step.
@@
Finally, the route is encoded in the bundle and the bundle is sent in the network. A bundle with an encapsulated route at a transmitting node (source or intermediate node) is subject to checks to verify whether this route can be used or not. The contact capacity and buffer managements during the forwarding phase in the benchmark are detailed hereafter.
@@ V-1 Contact-capacity management after the route search
First, the time at which the corresponding contact must be used for the bundle transmission is verified. In case of collision with another bundle, the transmission of the bundle having the highest ID is delayed (similar behavior to that of determining the ETO). The ID increases with the bundle generation time at the source. Two cases can occur:
  • If delaying the bundle does not cause any problem, the times of future hop uses are simply updated in the encapsulated route according to the delay caused by other bundles using the same contact.
@@ V-2 Buffer management after the route search
An alternative solution, which we are not considering, would be to continue trying to transmit via the fully booked node each time step until a valid route is found or a place becomes available in the buffer of the problematic node, whichever comes first. Note that this solution requires a lot of computing resources as the route search is repeated at each time step until a suitable solution is found.
@@ VI-A Environment
The CP is constructed for a 24-hour scenario starting from 8:55 pm on August 19, 2020. We assume that all the contacts have the same rate $R_{C_{j}}$.
@@
Figure 9: Snapshot of the satellite communication scenario with $N_{\text{sat}}=16$ at 11:18:08 pm on August 19, 2020. This corresponds to constellation 3 in Table I.
In the considered constellations, each orbital plane has $\frac{N_{\text{Sat}}}{N_{\text{p}}}$ satellites. We define the plane hops as the number of planes a communication between two satellites can cross (Table I). Depending on the communication parameters and the satellite dynamics, a satellite located in an orbital plane can communicate with satellites located in the same plane (zero plane hop, same RAAN), in adjacent planes (one plane hop, one $\Delta\Omega$ difference between RAANs of the satellites) or in more distant planes (more than one plane hops, meaning that there is no plane restriction in our case as the number of planes is limited).
The characteristics of the used constellations, also depending on the communication parameters provided in the following subsection, are given in Table I.
@@ VI-B Communication parameters
Example. In Figure 9, constellation $3$ is represented. The figure illustrates, in green, link closures between pairs of nodes: sat13-sat7, sat11-sat1, and sat14-GS 2, for example. These link closures happens at the time of the snapshot. The link between sat14 and GS 2 is at its limit: One time sample (of one second) later, this link no longer exists as these two nodes cease to respect the elevation angle constraint. The link closure is thus interrupted even though they remain in line-of-sight and in their communication range. In addition, sat4 and sat8 may be within communication range, but their communication is blocked by the Earth, preventing them from being in line-of-sight.
TABLE I: Constellation characteristics.
Constellation $N_{\text{p}}$ $N_{\text{sat}}$ Number of Mean contact Plane
@@ VI-C Other simulation parameters and assumptions
At each time step, one bundle of $100$ Bytes is generated at the source. The bundles are generated over a period of $500$ seconds, starting from the beginning of the scenario time, with equal inter-arrival times. In total $N_{\text{b}}$ bundles are generated and forwarded. This means that one bundle is generated each $\frac{500}{N_{\text{b}}}$ seconds (the bundle inter-arrival time decreases as $N_{\text{b}}$, the total number of bundles generated at source, increases).
@@ VI-D Simulation tools
We use MATLAB Satellite Communications Toolbox [22] to simulate the presented satellite communication scenario. We can then calculate the link closures between each pair of nodes as a function of time to obtain the CP used in the benchmark and in our solution, which are also implemented in MATLAB.
@@ VII-A Impact of contact-capacity management
Figure 10 shows the average time spent in the network for $N_{\text{sat}}=16$ (constellation $3$) as function of the number of bundles generated at source $N_{\text{b}}$. Three values of contact data rate $R_{C_{j}}$ are considered: $200$, $400$, and $800$ bps for all contacts. As expected, one can see that the time spent in the network increases with the number of bundles and with decreasing data rate. With low load (e.g., $N_{\text{b}}<300$ for $400$ bps), the proposed algorithm and the SR-CGR have similar performance. As $N_{\text{b}}$ increases, the gap between the curves increases, with less time spent in the network using the proposed method. This is due to the increased number of rerouting events during forwarding time when using SR-CGR. This number of rerouting events is shown in Figure 11 for the same $N_{\text{sat}}$ (constellation $3$) and $R_{C_{j}}$ values as function of $N_{\text{b}}$. With low load, there are only few overbooking events happening on the contacts. In this case, traditional SR-CGR has satisfactory performance. When the number of bundles increases, the overbooking phenomenon increases, making rerouting more likely. In case of such overbooking event, intermediate nodes try to overcome this issue by rerouting the bundle (from their location toward the destination). Such new routes are longer than the ones avoiding overbooking (computed using the proposed algorithm). Moreover, if no route is found from the node, the bundle has to be rerouted from the source (see the assumptions on the benchmark) also causing additional delay. The proposed algorithm is more efficient because the shortest route is always computed directly from the source node, always avoiding overbooking. Indeed, as expected, no rerouting was required with our proposed solution in the simulation due to contact capacity check failures.
@@
Moreover, although in some cases the time spent in the network with SR-CGR is similar to the proposed solution, the node complexity is significantly higher with the benchmark due to multiple checks and rerouting events.
@@
Then, Figure 12 illustrates the time spent in the network with $4$, $8$, and $16$ satellites (Constellations $1$, $2$, and $3$, respectively). Contact data rate is considered equal to $400$ bps. With four satellites (one satellite per orbital plane), the number of nodes is limited and they are more distant compared to constellations with higher number of satellites. Thus, the number of communication opportunities are limited, limiting the routing degrees of freedom. In this case, the proposed algorithm does not show improvement compared to the benchmark. When the number of satellites increases to $8$ or $16$, our algorithm presents improvement compared to the benchmark, in particular with high load ($N_{\text{b}}>250$).
Figure 12: Average time spent in the network ($R_{C_{j}}=400$ bps).
@@ VII-B Impact of buffer management
In this section, we evaluate the performance of the proposed buffer management. No contact capacity restriction is implemented, i.e., $R_{C_{j}}=\infty,\;\forall j$.
The average time spent by a bundle in the network as a function of the buffer size per node, expressed by the number of bundles a buffer can store, is given in Figure 13. We consider $N_{\text{sat}}=16$ (constellation $3$). The buffer storage size is identical for all intermediate nodes. The source node has infinite buffer size. The results are shown for $N_{\text{b}}=250$ and $N_{\text{b}}=100$ bundles generated during $500$ seconds. As expected, when the storage capacity of buffers increases, the time spent in the network is reduced. The gain is significant with the proposed solution, especially when the buffer sizes are limited. To reach the performance of infinite buffer capacity for $N_{\text{b}}=100$, our solution requires buffers that can store $10$ bundles, to be compared with $40$ bundles (almost half of $N_{\text{b}}$) to achieve this performance with the benchmark. The difference is even higher with $N_{\text{b}}=250$, where to reach the performance of infinite buffer the buffer sizes should be $40$ and $200$ bundles (small box in the figure) using our solution and the benchmark, respectively.
@@
Similarly to the capacity case, these differences are due to the number of rerouting events on intermediate nodes as shown in Figure 14, also for $N_{\text{sat}}=16$ (constellation $3$), when using SR-CGR. Moreover, as expected, no rerouting had to be performed with the proposed solution due to a buffer check failure.
@@
Finally, Figure 15 shows the impact of different constellations, namely $1$ and $3$, on the performance related to buffer constraints and for $N_{\text{b}}=250$. When the number of nodes is limited, each of them is more frequently used. The limited storage capacity then induces additional delay for both solutions: With $N_{\text{sat}}=~{}4$, the two solutions reach the infinite buffer performance at the same buffer capacity ($100$ bundles per node). Nevertheless, with that same constellation and with lower buffer size per node, our proposed solution manages to significantly reduce the average time a bundle spends in the network. Increasing the number of nodes in the constellation and using our proposed solution leads to very low delay even with low buffer capacity per node.
Figure 15: Average time spent in the network as function of buffer size for $N_{\text{b}}=250$.

### v2

@@ IV-C Booking resources for forwarding safety margin
In the proposed solution, only source nodes monitor buffer and contact capacity status. This centralized monitoring solution enables source nodes to prevent overflows and overbooking. However, unpredictable disruptions (such as antenna pointing errors or node power outages [35]) may force intermediate nodes to perform routing during the forwarding phase. Intermediate nodes lack access to current buffer and capacity status, risking overflows or overbooking with their routing decisions. To address this risk, we reserve portions of contact and buffer resources, keeping them unused by source nodes. These reserved portions, called safety margins, allow intermediate nodes to reroute bundles safely without causing overflows or overbooking.
The network’s disruption level determines the required resource allocation. This allocation requires careful tuning to optimize resource usage. The safety margin must balance two requirements: sufficient size to prevent bundle loss during intermediate node rerouting, while remaining minimal to preserve network efficiency.
These safety margins can be implemented as follows. First, regarding the contact-capacity management, two CPs are created. Contact splitting is performed for the slots that should be kept as margin. The safety margin is therefore removed from the original CP. The obtained CP is provided to and used by the source nodes. The safety-margin part (contacts removed from the original CP) is forwarded to intermediate nodes that may perform routing in unpredictable situations. In terms of buffer management, source nodes utilize a fraction of the buffer capacity, leaving the rest available for intermediate nodes. This solution therefore enhances robustness to potential unexpected events, despite the poor knowledge of the current network status by intermediate nodes.
@@ IV-D Information sharing in the network
The modified network information after contact splitting (except the bundle-specific CP modifications) and the modified forecast buffer tables (after allocating a route for a bundle) are to be shared with all source nodes performing a route search. This is similar to the sharing of the up-to-date CP with the source nodes in standard SR-CGR. The safety margin resources are shared with all nodes in the network at a low frequency.
@@ V Benchmark: source-routing CGR
As mentioned in the first part of the document, network issues are handled after the route search in standard SR-CGR. While re-transmission mechanisms do exist for long round-trip times communication, such as Licklider Transmission Protocol (LTP) [36], the re-transmission approach in case of failure in standard SR-CGR is not well described in the literature. As there are many manners to address these issues, we describe in details the assumptions made for the considered SR-CGR benchmark. The proposed algorithm is compared to this benchmark in Section VII, where the simulation results are presented.
To begin with, both in the benchmark and in our algorithm, when a bundle is being transmitted over a contact, it is assumed that the entire link is occupied for the time required to transmit the bundle. In other words, the whole link capacity is used (i.e., Time Division Multiplexing (TDM) approach). The initial CP is constructed at the source and is broadcasted to all nodes in the network.
Then, the source node performs a route search using Yen’s algorithm with Dijkstra searches. A list of the $k$ (Yen’s constant) shortest routes is generated. This list is kept until a change occurs in the CP, such as the end of a contact. In this case, the route list is discarded and a new list is generated, as in route-list pruning [28]. A list of candidate routes is constructed as explained in Section II-C3. When no candidate route is found in this list, the bundle is delayed and the route search is performed again after one time step.
@@
Finally, the route is encoded in the bundle and the bundle is sent in the network. A bundle with an encapsulated route at a transmitting node (source or intermediate node) is subject to checks to verify whether this route can be used or not. The contact capacity and buffer managements during the forwarding phase in this benchmark are detailed hereafter.
@@ V-1 Contact-capacity management after the route search
First, we verify the time at which the corresponding contact must be used for the bundle transmission. In the case of collision with another bundle, the transmission of the bundle having the highest ID is delayed (similar behavior to that of determining the ETO). The ID increases with the bundle generation time at the source. Two cases can occur:
  • If delaying the bundle does not cause any problem, the times of future hop uses are simply updated in the encapsulated route according to the delay caused by other bundles using the same contact. This update accounts for the delay caused by other bundles using the same contact.
@@ V-2 Buffer management after the route search
An alternative solution, which we are not considering, would be to continue trying to transmit via the fully booked node at each time step until a valid route is found or a place becomes available in the buffer of the problematic node, whichever comes first. Note that this solution requires a lot of computing resources as the route search is repeated at each time step until a suitable solution is found.
@@ VI-A Environment
The CP is constructed for a 24-hour scenario starting from 8:55 pm on August 19, 2020. We assume that all the contacts have the same rate $R_{\text{C}_{j}}$.
@@
Figure 7: Snapshot of the satellite communication scenario with $N_{\text{sat}}=16$ at 11:18:08 pm on August 19, 2020. This corresponds to constellation 3 in Table III.
In the considered constellations, each orbital plane has $\frac{N_{\text{Sat}}}{N_{\text{p}}}$ satellites. We define the plane hops as the number of planes a communication between two satellites can cross (Table III). Depending on the communication parameters and the satellite dynamics, a satellite located in an orbital plane can communicate with satellites located in the same plane (zero plane hop, same RAAN), in adjacent planes (one plane hop, one $\Delta\Omega$ difference between RAANs of the satellites) or in more distant planes (more than one plane hops, meaning that there is no plane restriction in our case as the number of planes is limited).
The characteristics of the used constellations, also depending on the communication parameters provided in the following subsection, are given in Table III.
@@ VI-B Communication parameters
Example. In Figure 7, constellation $3$ is represented. The figure illustrates, in green, link closures between pairs of nodes: sat13-sat7, sat11-sat1, and sat14-GS 2, for example. These link closures happen at the time of the snapshot. The link between sat14 and GS 2 is at its limit: One time sample (of one second) later, this link no longer exists as these two nodes cease to respect the elevation angle constraint. The link closure is thus interrupted even though they remain in line-of-sight and in their communication range. In addition, sat4 and sat8 may be within communication range, but their communication is blocked by the Earth, preventing them from being in line-of-sight.
TABLE III: Constellation characteristics.
Constellation $N_{\text{p}}$ $N_{\text{sat}}$ Initial number of Mean contact Plane
@@ VI-C Other simulation parameters and assumptions
At each time step, one bundle of $100$ Bytes is generated at the source. The bundles are generated over a period of $2000$ seconds, starting from the beginning of the scenario time, with equal inter-arrival times. In total $N_{\text{b}}$ bundles are generated and forwarded. This means that one bundle is generated each $\frac{2000}{N_{\text{b}}}$ seconds (the bundle inter-arrival time decreases as $N_{\text{b}}$, the total number of bundles generated at source, increases).
@@ VI-D Simulation tools
We use MATLAB Satellite Communications Toolbox [38] to simulate the presented satellite communication scenario. We can then calculate the link closures between each pair of nodes as a function of time to obtain the CP used in the benchmark and in our solution, also implemented in MATLAB.
@@ VII-A Impact of contact-capacity management
Figure 8 shows the average time spent in the network for $N_{\text{sat}}=16$ (constellation $3$) as function of the number of bundles generated at source $N_{\text{b}}$. Three values of contact data rate $R_{\text{C}_{j}}$ are considered: $200$, $400$, and $800$ bps for all contacts. As expected, one can see that the time spent in the network increases with the number of bundles and with decreasing data rate. With low load (e.g., $N_{\text{b}}<400$ for $400$ bps), the proposed algorithm and the SR-CGR have similar performance. As $N_{\text{b}}$ increases, the gap between the curves increases, with less time spent in the network when using our proposed method. This is due to the increased number of rerouting events during forwarding time when using SR-CGR. This number of rerouting events is shown in Figure 9 for the same $N_{\text{sat}}$ (constellation $3$) and $R_{\text{C}_{j}}$ values as function of $N_{\text{b}}$. With low load, there are only few overbooking events happening on the contacts. In this case, traditional SR-CGR has satisfactory performance. When the number of bundles increases, the overbooking phenomenon increases, making rerouting more likely. In case of such overbooking events, intermediate nodes try to overcome this issue by rerouting the bundle from their location toward the destination. Such new routes are longer than the ones avoiding overbooking, computed using the proposed algorithm. Moreover, if no route is found from the node, the bundle has to be rerouted from the source (see the assumptions on the benchmark) also causing additional delay. The proposed algorithm is more efficient because it computes the shortest route directly from the source node while avoiding overbooking. Indeed, as expected, our proposed solution required no rerouting due to contact capacity check failures during the simulation.
@@
Moreover, although in some cases the time spent in the network with SR-CGR is similar to the proposed solution, the intermediate node complexity is significantly higher with the benchmark due to multiple checks and rerouting events.
@@
Then, Figure 10 illustrates the time spent in the network with $4$, $8$, and $16$ satellites (Constellations $1$, $2$, and $3$, respectively). The contact data rate is $400$ bps. With four satellites (one satellite per orbital plane), the number of nodes is limited and they are more distant compared to constellations with higher numbers of satellites. Thus, the number of communication opportunities is limited, which constrains the routing degrees of freedom. In this case, the proposed algorithm shows a small improvement compared to the benchmark as seen in Figure 10(a). When the number of satellites increases to $8$ or $16$, our algorithm shows greater improvements compared to the benchmark, particularly with high load ($N_{\text{b}}>1000$ in figures 10(b) and 10(c)).
(a) $N_{\text{sat}}=4$.
(b) $N_{\text{sat}}=8$.
(c) $N_{\text{sat}}=16$.
Figure 10: Average time spent in the network ($R_{\text{C}_{j}}=400$ bps).
Figure 11 illustrates the evolution of the CP size over time, to assess the complexity induced by the proposed algorithm. The Y-axis shows the number of bundles in the CP normalized by its initial size, which is of $3478$ contacts for Constellation $3$ (see Table III). The baseline curve, without capacity management, depicts the evolution of the CP size over time as contacts expire and are removed from the CP. Newly appearing contacts are not counted in this scenario. The other curves represent the evolution of the CP size with our capacity management solution. In this case, splitting contacts can either increase the CP size or leave it unchanged if contacts are shortened. Expired contacts are still removed from the CP. We observe that our capacity management has a negligible impact on the CP size and route search complexity under low load conditions. For instance, with $N_{\text{b}}=200$, the maximum difference between the baseline and our split CP size is $1\%$ and $0.3\%$ for $R_{\text{C}_{j}}=200$ and $800$ bps, respectively. Under high load conditions, the difference becomes more significant. For example, with $N_{\text{b}}=2000$, the increased frequency of bundle generation and routing forces the use of late contacts, as early contacts become fully booked. Consequently, the CP size remains elevated until both split and original contacts expire. The highest difference with $N_{\text{b}}=2000$, observed for $R_{\text{C}_{j}}=200$ at $t=1920$, is approximately $10\%$ compared to the baseline.
Figure 11: Impact of capacity management on the CP size, $N_{\text{sat}}=16$.
@@ VII-B Impact of buffer management
In this section, we evaluate the performance of the proposed buffer management solution. No contact capacity restriction is implemented, i.e., $R_{\text{C}_{j}}=\infty,\;\forall j$.
The average time spent by a bundle in the network as a function of the buffer size per node, expressed by the number of bundles a buffer can store, is given in Figure 12. We consider $N_{\text{sat}}=16$ (constellation $3$). The buffer storage size is identical for all intermediate nodes. The source node has an infinite buffer size. The results are shown for $N_{\text{b}}=1000$ (Figure 12(a)) and $N_{\text{b}}=2000$ (Figure 12(b)) bundles generated over $2000$ seconds. As expected, when the storage capacity of the buffers increases, the time spent in the network is reduced. The gain is significant with the proposed solution, especially when the buffer sizes are limited. To reach a performance comparable to infinite buffer capacity for $N_{\text{b}}=1000$, our solution requires buffers that can store $100$ bundles, to be compared with $200$ bundles to achieve this performance with the benchmark. The performance gap becomes even more pronounced with $N_{\text{b}}=2000$: our solution achieves infinite-buffer performance with only $100$ bundles of buffer space, while the benchmark requires more than $300$ bundles to reach the same performance level. Even with low buffer capacity, our buffer management performs comparably to an infinite buffer capacity case, unlike the benchmark. For example, with a buffer size of $50$ bundles, our buffer management results in a $17\%$ increase compared to an infinite buffer capacity, whereas the benchmark shows a $2000\%$ increase for $N_{\text{b}}=2000$.
(a) $N_{\text{b}}=1000$.
(b) $N_{\text{b}}=2000$.
@@
Similarly to the capacity case, these differences are due to the number of rerouting events on intermediate nodes as shown in Figure 13, also for $N_{\text{sat}}=16$ (constellation $3$), when using SR-CGR. As expected, no rerouting had to be performed with the proposed solution.
@@
Figure 14 illustrates the performance impact of constellations with different sizes ($1$, $2$, and $3$) under buffer constraints, using $N_{\text{b}}=400$. In networks with fewer nodes, each node experiences higher utilization. With $N_{\text{sat}}=4$, the limited number of communication opportunities results in similar performance between both solutions, as the additional delay from limited storage capacity affects both approaches equally. Increasing the number of satellites to $8$ highlights the effectiveness of our solution, which results in less time spent in the network than the benchmark. Additionally, when the constellation is expanded to $16$ satellites, our proposed solution achieves very low delay, even with low buffer capacity per node.
Figure 14: Average time spent in the network as function of buffer size for $N_{\text{b}}=400$.
In Figure 15, we represent the evolution of the CP size, normalized by its initial size ($3478$, see Table III), as function of time. The size of the CP is reported each second. Without splitting for buffer management, the size of the CP decreases with time as we only consider that the expired contacts are removed from the CP as time advances. When considering buffer management, the splits are temporarily performed at bundle generation/routing time. With $N_{\text{b}}=2000$, one bundle is generated each second. When the buffer size is $50$ bundles, the temporal splitting impacts the CP size at almost all the time samples (also equal to the bundle generation time in this case). The impact of buffer management in this case is negligible, where the maximum is an increase of $0.43\%$ in the CP size at $1992$ seconds. When the buffer size increases to $500$ bundles for the same $N_{\text{b}}$, the CP size is not impacted most of the times due to high buffer capacity. It is nevertheless impacted between $1140$ and $1720$ seconds, where the maximum temporal increase is also negligible ($0.11\%$). We observe similar performance with $N_{\text{b}}=200$ and a buffer size of $50$ bundles. Here, one bundle is generated and routed every $10$ seconds, causing a temporary increase in the CP size, if it occurs, only during packet generation and routing. Outside of these times, the CP size remains equal to its original size, without splitting.
Figure 15: Impact of buffer management on the CP size.